The Roth family has liquid assets of 10,000, use assets of 150,000 and investment assets of 34,000. They also have liabilities totaling $108,000. If their total liabilities increase by $7,000 over the next year, by how much would the Roth family need to increase their assets in order to have a net worth of $100,000?
1
Expert's answer
2020-10-18T17:49:47-0400

Total assets "= \$10,000 + \$150,000 + \$34,000 = \$ 194,000,"

Total present liabilities = $108,000,

After increasing liabilities by $ 7,000,

New liabilities "= \$108,000 + \$ 7,000 = \$ 115,000,"

Let family need to increase their assets by x dollars in order to have a net worth of $100,000,

∵ Net worth = Assets - Liabilities

"100,000 = 194,000 + x - 115,000"

"100,000 = 79000 + x"

⇒ "x = 100,000 - 79000"

"= \$21000"